# Improving Sequential Recommendations with LLMs

### v1

@@ 4.4. Fine-tuning Task: Rank Next Item Recommendations
The ranking task enables the fine-tuned model to draw its attention to the ranking of recommendations. By fine-tuning a model to rank recommendations, we aim to investigate the efficacy in the ranking task of an LLM adapted to a particular domain compared to the more open fine-tuning task specifications of Sections 4.14.2, and 4.3.
In this variant, visualized in Figure 6, we fine-tune the models to learn a ranking task where we pass in the prompt a session of items except the last one and a set of recommendations including the ground truth, that is the last item of the session. We take the recommendations from an existing recommendation model depending on the dataset [^15] We chose LLMSeqSim for Amazon Beauty. similarly to Section 4.2, and shuffle them. Then, in the prompt completion we provide the recommendations in the order produced by the model and ensure that the ground truth item is at the top of the list, putting or moving it there as necessary. At prediction time, we ask the fine-tuned model to re-rank the recommendations in the prompt based on what would be the best completion for a prompt session. Similarly to the ranking problem formulation, no post-processing of recommendations is required, since the model chooses between options provided in the prompt.
@@ 7.1. Experimental setup
The experimental setup consists of the datasets and their pre-processing (Section 7.1.1), the metrics used to evaluate experimental results (Section 7.1.2), the models that were included in the experiments (Section 7.1.3), the hyperparameter tuning process for executing models with their best configurations (Section 7.1.4), and precise information regarding which models were executed on which datasets, including exceptions (Section 7.1.5).
@@ 7.1.1. Datasets and Data Splitting
We use the public Amazon Beauty (He and McAuley, 2016) dataset [^18] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html, Section Per-category file., a novel, real-world proprietary dataset from Delivery Hero, and the public Steam dataset [^19] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data; Version 2: Review Data and Item Metadata. from the gaming domain. The Beauty dataset contains product reviews and ratings from Amazon. In line with prior research (Anelli et al., 2022), we pre-processed the dataset so that there are at least five interactions per user and item ($\mathop{p-\mathrm{core}}=5$). The Delivery Hero dataset contains anonymous QCommerce sessions for dark store and local shop orders. [^20] QCommerce is a segment of e-Commerce focusing on fast delivery times on the last mile. To better simulate a real-world setting, we did not pre-process this dataset, except that we removed sessions with only one interaction. In the Steam dataset we applied $\mathop{p-\mathrm{core}}=5$ pre-processing and removed interactions that contain items with insufficient metadata (title, genres, and tags). Dataset statistics are provided in Table 2.
We selected a diverse set of datasets to ensure that our findings are not limited to datasets with particular characteristics. The datasets are of different nature, originate from different domains, and have different features. In the left part of Figure 11, we visualize the popularity of items in the sessions. After counting the number of occurrences of each item, we rank them in descending order (the most popular being the first) and then plot the log of the frequency on the log of the rank. The Delivery Hero dataset has a tail of items that appear 1 to 4 times, as it is not $\mathop{p-\mathrm{core}}=5$. In the right part of the figure, we show the log of the frequency of the sessions’ length. The Delivery Hero dataset has shorter sessions than the Steam dataset, and fewer longer sessions (those with more than 40 items) than either of the other two datasets.
@@ 7.1.3. Models
We include both session-based algorithms of different families, GRU4Rec (Hidasi et al., 2016), and SKNN (Jannach and Ludewig, 2017), as well as two state-of-the-art sequential models, BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018). We tested all variants of the SKNN nearest-neighbor method proposed in (Ludewig and Jannach, 2018) and report the results in the online material. In addition, we include the four LLM-based approaches proposed in Sections 34,  5, and 6. Finally, we include a popularity-based baseline (MostPopular) in the experiments.
7.1.4. Hyperparameter Tuning
We systematically tuned all models (except LLMSeqPrompt) on three validation folds with the Tree Parzen Estimator (TPE) sampler (Bergstra et al., 2011), and the total NDCG@20 across the folds as the optimization goal. We chose to optimize for 20 recommendations to provide enough room for the models to manifest their performance differences, although top-10 and top-20 metrics are usually closely related. In the beginning, we explore 40 random configurations, which helps avoid local minima, and then continue by evaluating the suggestions by the TPE sampler. We let the hyperparameter search run for 72 hours, but started multiple hyperparameter search experiments in parallel for each model (two for vanilla neural models and four for LLM2Sequential models) in order to minimize the duration of experiments. We also allow for early stopping of the hyperparameter search if the optimization objective has not been improved for 100 trials. Finally, we prune a trial if its objective value (i.e., NDCG@20) belongs to the bottom 20% of all tested trials up to that point, allowing us to accelerate the hyperparameter search process. We allow pruning only after ten different configurations of a model have been tried.
7.1.5. Model Executions on Datasets
In principle, we ran all models on all datasets except for fine-tuning models due to budget and time limitations. Specifically, we applied LLMSeqPromptGenList, LLMSeqPromptClass, and LLMSeqPromptRank only on the Beauty dataset. In addition, we only fine-tuned OpenAI ada on LLMSeqPromptGenItem because it is no longer available for fine-tuning.
@@ 7.2.2. Beyond accuracy metrics
We make the following observations for coverage, serendipity, and novelty. The LLMSeqSim model consistently leads to the best catalog coverage and novelty. This is not too surprising, given the nature of the approach, which is solely based on embedding similarities. Unlike other methods that use collaborative signals, i.e., past user-item interactions, the general popularity of an item in terms of the amount of observed past interactions does not play a role in LLMSeqSim, neither directly nor implicitly. Thus, the model has no tendency to concentrate the recommendations on a certain subset of (popular) items. [^22] We recall that the used novelty measure is based on the popularity of the items in the recommendations. Finally, the serendipity results are aligned with the accuracy measures across the datasets. This generally confirms the value of personalizing the recommendations to individual user preferences, compared to recommending mostly popular items to everyone. [^23] We iterate that our serendipity measure counts the fraction of correctly recommended items that would not be recommended by a popularity-based approach.
@@ 7.3.3. Fine-tuned Models (LLMSeqPrompt)
In addition, the LLMSeqPromptGenItem task has been used to fine-tune both ada and GPT 3.5 turbo OpenAI models on the Beauty and Delivery Hero datasets. The results are conflicting. On the one hand, we observe a notable performance increase of 23% in terms of NDCG@20 in favor of GPT on Beauty, but, surprisingly, on Delivery Hero ada performed significantly better than GPT for the same task achieving 18% higher NDCG@20. Given that the comparison scope is narrow, we can only conclude that the choice of fine-tune model depends on the dataset and that the evolution of models does not necessarily lead to higher performance.
With respect to the temperature and top_p hyperparameters we observe that in all setups a high temperature results in lower performance as it allows for more creativity at the cost of item hallucinations. For LLMSeqPromptGenItem and LLMSeqPromptGenList a low but non-zero temperature leads to the best performance whereas LLMSeqPromptClass and LLMSeqPromptRank perform best with temperature set to zero. A possible explanation for this is that the latter variants do not require any creativity because the recommendation options are contained in the provided prompt.
@@
Finally, to explore the learning capabilities of fine-tuned models we compare their performance to the corresponding base models. We observe that in all setups the fine-tuned models significantly outperform their non fine-tuned counterparts. For LLMSeqPromptGenItem and LLMSeqPromptGenList variants the model generates recommendations that are not part of the prompt. For those, the base model has a much higher rate of hallucinations compared to fine-tuned models as Table 7 shows. Additionally, the hallucinations produced by the fine-tuned models exhibit a higher embedding similarity to the catalog items compared to the base model. These observations indicate that fine-tuned models recall and prioritize concepts that they processed during their fine-tuning compared to the base model, which is not given the same opportunity. Therefore, the base model has to provide recommendations solely based on the information in the prompt.
For variants LLMSeqPromptClass and LLMSeqPromptRank the recommendation options are explicitly part of the context such that the model only has to select and/or re-rank the given item names. Thus, hallucinations are uncommon in those variants. Still, the performance advantage of the fine-tuned models over the base models persists. This indicates that the fine-tuned models learn the underlying classification and ranking tasks better. This is an expected finding that is widely acknowledged by literature and practice (, 2023; Touvron et al., 2023).
All in all, the fine-tuned models perform better for each recommendation task with less hallucinations and closer semantic proximity to the actual options compared to the base models. This finding clearly shows that fine-tuning can enable a model to learn not only a task, but also concepts of a domain to some extent and contradicts existing findings and common belief that fine-tuning is only applicable for learning form rather than facts (Mitra et al., 2023; Kirstain et al., 2022; Ovadia et al., 2024; Lewis et al., 2021; Wang et al., 2023; Kandpal et al., 2023; Luo et al., 2023; Chen et al., 2020; Kirkpatrick et al., 2017; Berglund et al., 2023).
@@ 8. Challenges, Limitations, & Conclusions
In terms of research limitations, our work so far is based on three datasets from two domains (e-commerce, games). As the datasets used in our research are markedly different in terms of various characteristics (e.g., sparsity, size, available meta-data), we are confident that our findings generalize beyond the conducted experiments. Nonetheless, further experiments are needed to validate the benefits of leveraging LLMs for sequential recommendation in other domains, e.g., in music or news recommendation.

### v2

@@ 4.4. Fine-tuning Task: Rank Next Item Recommendations
In this final variant, visualized in Figure 6, we fine-tune the models to learn a ranking task. Specifically, we create a prompt that consists of (a) a session of items except the last one, and (b) a set of recommendations including the ground truth, that is the last item of the session. We take the set of recommendations from an existing recommendation model depending on the dataset [^15] We chose LLMSeqSim for Amazon Beauty. similarly to Section 4.2, and shuffle them. Then, in the prompt completion we provide the recommendations in the order produced by the model and ensure that the ground truth item is at the top of the list, putting or moving it there as necessary. At prediction time, we ask the fine-tuned model to re-rank the recommendations in the prompt based on what would be the best completion for a prompt session.
@@ 5.2. LLM-enhanced Sequential Models
In terms of technical aspects, we note that the item embeddings in our approach are trainable and they are computed based on item metadata of the domain. For the Beauty and the Delivery Hero datasets, the metadata used to compute the embeddings are the names of the products. In the Steam dataset, the items of the domain are games. Typically, the names of the games are not closely linked with the concept of a game. Thus, we concatenate the names of the games with tags that accompany and characterize each game. The IDs of the items are not included to the input of the embedding computation.
In the overall process, the collaborative information thus enters the LLM2Sequential model as a sequence of IDs representing the items of the domain that appear in the sequence. In the embedding layer, the sequence of IDs becomes a sequence of embeddings in the following manner. First, the item embedding of each item in the sequence is retrieved from the embedding matrix based on the ID of the item. The item embedding is the embedding previously received by an LLM based on metadata of the item provided as input to the LLM. Second, each item embedding is summed with the positional embedding of the item’s position in the sequence to form the final embedding representation. Then, the embeddings are transformed while passing through layers of the model in the feed-forward pass where the collaborative and textual information are fused in the multilayer perceptron of the feed-forward network. Finally, through backpropagation of the gradients that minimize the cost, the weights of each layer are updated leading to the update of the embeddings in the embedding layer as well.
Neighborhood-based Models
@@ 7.1. Experimental setup
The experimental setup consists of the datasets and their pre-processing (Section 7.1.1), the metrics used to evaluate experimental results (Section 7.1.2), the models that were included in the experiments (Section 7.1.3), the hyperparameter tuning process for executing models with their best configurations (Section 7.1.4), the experimental setup of S3-Rec (Section 7.1.5), TALLRec (Section 7.1.6), and LlamaSeqPromptGenItem (Section 7.1.7) as well as precise information regarding which models were executed on which datasets, including exceptions (Section 7.1.8).
@@ 7.1.1. Datasets and Data Splitting
[…]
@@ 7.1.3. Models
We include both session-based algorithms of different families, GRU4Rec (Hidasi et al., 2016), and SKNN (Jannach and Ludewig, 2017), as well as three state-of-the-art sequential models, BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), and S3-Rec (Zhou et al., 2020). Notably, S3-Rec introduces a custom pre-training and fine-tuning method that allows a meaningful comparison to our proposed LLM-based approaches, especially with respect to the LLM embedding initialization of neural models (Section 5) and finetuning of LLMs (Section 4). We also provide a comparison with a fine-tuned open-source Llama (Touvron et al., 2023) LLM. In addition, we evaluate TALLRec (Bao et al., 2023), a state of the art LLM-based approach for sequential recommendations that structures the recommendations as instructions and tunes an LLM through an instruction-tuning process. Both Llama and TALLRec are closely related to our proposed LLM fine-tuning approaches (Section 4). Additionally, we initialize neural models with embeddings retrieved from a pre-trained BERT (Devlin et al., 2019) model to compare against our proposed initialization approach using embeddings retrieved from OpenAI GPT and Google PaLM (Section 5). We refer to those models as BERT2BERT4Rec, BERT2SASRec, and BERT2GRU4Rec. Furthermore, we tested all variants of the SKNN nearest-neighbor method proposed in (Ludewig and Jannach, 2018) and report the results of the best SKNN variant in this paper. The complete results can be found in the online material. Moreover, we include the four LLM-based approaches proposed in Sections 34,  5, and 6. Finally, we include a popularity-based baseline (MostPopular) in the experiments. [^22] The consideration of additional traditional baselines and most recently published sequential models is left for future work.
@@
7.1.4. Hyperparameter Tuning
We systematically tuned all models (except LLMSeqPrompt, LlamaSeqPromptGenItem, S3-Rec, and TALLRec, which we discuss separately) on three validation folds with the Tree Parzen Estimator (TPE) sampler (Bergstra et al., 2011), and the total NDCG@20 across the folds as the optimization goal. We chose to optimize for 20 recommendations to provide enough room for the models to manifest their performance differences, although top-10 and top-20 metrics are usually closely related. In the beginning, we explore 40 random configurations, which helps avoid local minima, and then continue by evaluating the suggestions by the TPE sampler. We let the hyperparameter search run for 72 hours, but started multiple hyperparameter search experiments in parallel for each model (two for vanilla neural models and four for LLM2Sequential models) in order to minimize the duration of experiments. We also allow for early stopping of the hyperparameter search if the optimization objective has not been improved for 100 trials. Finally, we prune a trial if its objective value (i.e., NDCG@20) belongs to the bottom 20% of all tested trials up to that point, allowing us to accelerate the hyperparameter search process. We allow pruning only after ten different configurations of a model have been tried.
7.1.5. Training and Evaluation of S3-Rec
The S3-Rec model is different in nature from other models, because it is a pre-training based approach. Nonetheless, we were interested how this particular model would fare in a comparison with the other models in our study. In terms of datasets, we included experiments only for the Beauty and Steam datasets, as these include metadata about item attributes that are required to pre-train S3-Rec. Using the original code of S3-Rec provided by the authors, we pre-trained S3-Rec for 10 epochs, fine-tuned it for 200 epochs with early stopping enabled. As for the other models in our comparison, we used the NDCG and the hit rate as evaluation measures. We note that we determined the metric values using the evaluation protocol from the original paper. S3-Rec’s evaluation protocol involves using each session’s second-to-last item for validation during fine-tuning. The last item of each session is used for testing. We note that the evaluation protocol (and code) used for the other models, which is more common in the literature, is slightly different and more challenging for the algorithms, as we hide entire sessions from the validation and training data. As a result, the results reported for the S3-Rec model are expected to be higher than what we would observe if the exact same protocol would be applied. As our results will show, however, the accuracy results obtained with S3-Rec fall behind the other models, even though more information was available than for the other models.
7.1.6. Training and Evaluation of TALLRec
For TALLRec, we present experiments only for the Beauty dataset because the model would require many weeks of execution on the Delivery Hero and Steam datasets. Regarding the model’s training, we used the same preprocessing and train-validation-test split for Beauty as for LLMSeqPromptGenItem (Section 4.1). To make TALLRec comparable to our work, we converted TALLRec’s binary classification problem into a recommendation generation problem. We also used the same system prompt, user prompt template, and assistant prompt template as in LLMSeqPromptGenItem but we adapted it to TALLRec’s format. [^23] https://github.com/SAI990323/TALLRec/blob/main/finetune_rec.py#L301 We then trained TALLRec on an EC2 g5.12xlarge instance with 4x24GB A10G GPUs using the default hyperparameter setup provided by the authors. [^24] https://github.com/SAI990323/TALLRec/blob/main/shell/instruct_7B.sh
For the evaluation, we used the same protocol as for LLMSeqPromptGenItem. After resolving out-of-memory issues, we asked the model to produce top-20 recommendations for each test case by repeating each test prompt 20 times. Then, we mapped each recommendation of the model that is not in the product catalog to the closest catalog product via OpenAI embeddings. Finally, we computed the evaluation metrics that we report in the paper.
7.1.7. Fine-tuning and Evaluation of LlamaSeqPromptGenItem
We fine-tuned LlamaSeqPromptGenItem, specifically meta-textgeneration-llama-3-8b, on the Beauty and Steam datasets using the LoRA approach for five epochs with 0.0001 learning rate, 32 alpha, 0.05 dropout, and we set the r parameter to 8. In addition, we used 20% of the dataset as validation set. Resource constraints did not allow us to conduct experiments on the Delivery Hero dataset. On the same grounds, we fixed the temperature of the model to 1 and the top_p parameter to 0.25, which perform best in practice according to our experience.
7.1.8. Model Executions on Datasets
In principle, we ran all models on all datasets except for fine-tuning models due to budget and time limitations. Specifically, we applied LLMSeqPromptGenList, LLMSeqPromptClass, LLMSeqPromptRank, and TALLRec only on the Beauty dataset, and S3-Rec and LlamaSeqPromptGenItem only on the Beauty and Steam datasets. In addition, we only fine-tuned OpenAI ada on LLMSeqPromptGenItem because it is no longer available for fine-tuning.
@@ 7.2.1. Accuracy metrics
In line with our observations regarding LLM embeddings, BERT embeddings also tend to increase the performance of neural recommendation models, but not consistently for all datasets and all models. In most of the cases, LLM embeddings significantly outperform BERT embeddings except for one case, for the DH dataset, where BERT embeddings elevated GRU4Rec’s performance significantly stronger than LLM embeddings. In summary, the experiments reveal a real potential for leveraging the semantics of a dataset’s item descriptions to uplift the performance of the recommendation models. This potential however depends on the dataset and different models capture the semantics in different ways and to different extents. Most of the time, however, embedding initialization leads to significant performance gains especially when LLM embeddings are used.
TALLRec, S3-Rec, and LlamaSeqPromptGenItem performed average in the evaluation. Specifically, they trail behind the neural models for the Steam dataset and also behind the LLM generation-based approaches for the Beauty dataset.
@@ 7.2.2. Beyond accuracy metrics
We make the following observations for coverage, serendipity, and novelty. The LLMSeqSim model consistently leads to the best catalog coverage and novelty except for Novelty@20 in the Beauty dataset where TALLRec performs better. The performance of LLMSeqSim in terms of catalog coverage and novelty is not too surprising, given the nature of the approach, which is solely based on embedding similarities. Unlike other methods that use collaborative signals, i.e., past user-item interactions, the general popularity of an item in terms of the amount of observed past interactions does not play a role in LLMSeqSim, neither directly nor implicitly. Thus, the model has no tendency to concentrate the recommendations on a certain subset of (popular) items. [^25] We recall that the used novelty measure is based on the popularity of the items in the recommendations. Finally, the serendipity results are aligned with the accuracy measures across the datasets. This generally confirms the value of personalizing the recommendations to individual user preferences, compared to recommending mostly popular items to everyone. [^26] We iterate that our serendipity measure counts the fraction of correctly recommended items that would not be recommended by a popularity-based approach.
@@ 7.3.3. Fine-tuned Models (LLMSeqPrompt)
In addition, the LLMSeqPromptGenItem task has been used to fine-tune both ada and GPT 3.5 turbo OpenAI models on the Beauty and Delivery Hero datasets. The results are inconclusive. On the one hand, we observe a notable performance increase of 23% in terms of NDCG@20 in favor of GPT on Beauty, but, surprisingly, on Delivery Hero ada performed significantly better than GPT for the same task achieving 18% higher NDCG@20. Given that the comparison scope is narrow, we can only conclude that the choice of fine-tune model depends on the dataset and that the evolution of models does not necessarily lead to higher performance.
With respect to the temperature and top_p hyperparameters, we observe that in all setups a high temperature results in lower performance as it allows for more creativity at the cost of item hallucinations. For LLMSeqPromptGenItem and LLMSeqPromptGenList a low but non-zero temperature leads to the best performance whereas LLMSeqPromptClass and LLMSeqPromptRank perform best with temperature set to zero. A possible explanation for this is that the latter variants do not require any creativity because the recommendation options are contained in the provided prompt.
@@
Finally, to explore the learning capabilities of fine-tuned models, we compare their performance to the corresponding base models. We observe that in all setups the fine-tuned models significantly outperform their non fine-tuned counterparts. For the LLMSeqPromptGenItem and LLMSeqPromptGenList variants, the model generates recommendations that are not part of the prompt. For those, the base model has a much higher rate of hallucinations compared to fine-tuned models as Table 7 shows. Additionally, the hallucinations produced by the fine-tuned models exhibit a higher embedding similarity to the catalog items compared to the base model. These observations indicate that fine-tuned models recall and prioritize concepts that they processed during their fine-tuning compared to the base model, which is not given the same opportunity. Therefore, the base model has to provide recommendations solely based on the information in the prompt.
For the variants LLMSeqPromptClass and LLMSeqPromptRank, the recommendation options are explicitly part of the context such that the model only has to select and/or re-rank the given item names. Thus, hallucinations are uncommon in those variants. Still, the performance advantage of the fine-tuned models over the base models persists. This indicates that the fine-tuned models learn the underlying classification and ranking tasks better. This is an expected finding that is widely acknowledged by literature and practice ( , 2023; Touvron et al., 2023).
All in all, the fine-tuned models perform better for each recommendation task with fewer hallucinations and closer semantic proximity to the actual options compared to the base models. This finding clearly shows that fine-tuning can enable a model to learn not only how to perform a task, but also concepts of a domain to some extent. Thus, it contradicts existing findings and common belief that fine-tuning is only applicable for learning form (instructions) rather than facts (Mitra et al., 2023; Kirstain et al., 2022; Ovadia et al., 2024; Lewis et al., 2021; Wang et al., 2023; Kandpal et al., 2023; Luo et al., 2023; Chen et al., 2020; Kirkpatrick et al., 2017; Berglund et al., 2023).
@@ 8. Challenges, Limitations, & Conclusions
In terms of research limitations, our work so far is based on three datasets from two domains (e-commerce, games). The datasets used in our research are however markedly different in terms of various characteristics (e.g., sparsity, size, available metadata, recurrence patterns, frequency of use) and in terms of their nature (e.g., one long sequence of product reviews per user for Amazon Beauty vs. a collection of supermarket purchases per user, each consisting of a sequence of products, in the Delivery Hero dataset). Therefore, we are confident that our findings generalize beyond the conducted experiments. Nonetheless, further experiments are needed to validate the benefits of leveraging LLMs for sequential recommendation in other domains, e.g., in music or news recommendation. In addition, in our current work we have not specifically analyzed the potential of LLM-enhanced sequential recommendations for cold-start situations, where very limited information is available about the preferences of individual users. Also, as another area for future works, it would be interesting to study the use of LLMs for cross-domain recommendation settings. Finally, future work might include the consideration of alternative LLMs or larger versions of existing LLMs, as these models continue to become more powerful and larger in terms of their parameters. One important research question in that context is if larger model sizes will have a significant impact on recommendation accuracy for the different approaches studied in this paper, see also the analyses in (Zhang et al., 2024).